# DDBot: Differentiable Physics-based Digging Robot for Unknown Granular Materials

### v3

@@ III-E Stopping gradients from exploding
A critical issue that destabilises or even disables gradient-based optimisation with differentiable physics is the numerical instability of gradient computation caused by long task horizons and the large particle number. As mentioned in existing works, gradients tend to explode or vanish as the length of the computation chain (number of actions/simulation steps) or the number of particles increases [8, 2]. This effect can be predicted based on Eq. 35, where $\partial o_{t+1}/\partial a_{t}$ depends heavily on the number of simulation steps $T$, substeps $N$, and particles $|o_{t+1}|$. It increases drastically (explodes) when there are too many sub-gradients that are greater than 11 ($\partial f_{sub}/\partial\boldsymbol{x}_{p}>1$), and vice versa (vanishes). This problem is exacerbated when the physical model becomes more complex, preventing previous works from using efficient first-order optimisation methods on tasks with long horizons or large numbers of particles [8, 9, 2, 3].
@@ IV-A Real-world platform
We set up a UR5e robotic arm to perform digging manipulations in the real world. As shown in Figure 3, the robot is equipped with a 3D-printed shovel as the end-effector. We place a $0.28\times 0.28$ m2 wooden box with $0.11$ metre depth to contain sand or soil for manipulation next to the robot. In the simulation, we use a box of the same size and the 3D model of the shovel as the simulated end-effector. We fill the container with the material, which could be sand or solid, to a height of $0.07$ metres. A Zivid Medium One+ high-resolution camera [37] is used to collect real-world surface point clouds.
@@ IV-B Soil/sand system identification
Before running skill learning and parameter optimisation with the simulation dynamics, an important task is to ensure that the simulated soil/sand dynamics match the real ones as closely as possible. To achieve this, we follow the Differentiable Physics-based System Identification (DPSI) framework introduced in work [5], which can identify physics parameters with only one datapoint. In simple terms, the DPSI framework adjusts the key parameters of a physics model using gradients backpropagated from a loss function that compares the simulated and real point clouds of the same manipulation motion.
@@
The ranges of the physics parameters are defined according to an open-source structural engineering dataset [38]. They are summarised together with the gradient descent step sizes in Table II. The RMSprop algorithm with $\beta=0.9$ [15] is used for gradient-based optimisation in this work.
@@
TABLE II: Stepsizes and the ranges of values for the physics parameters: Young’s modulus $E$, Poisson’s ratio $\nu$, material density $\rho$, and sand friction angle $\phi_{f}$, selected based on an open-source structural engineering dataset [38].
The feasibility of DPSI for physics parameter identification has been demonstrated with playdough-like plastic objects, which are governed by the fixed-corotated elastic energy model and the von Mises plasticity model [5]. In this paper, we cope with granular materials that are modelled by the St. Venant-Kirchoff (SVK) elastic energy and the Drucker-Prager (DP) plastic yield criterion [13].
@@ IV-C Digging tasks
After the simulation is aligned with the real materials, we examine the feasibility of gradient-based skill optimisation. We manually use the shovel to create $3$ digging target shapes with both soil and sand. The optical camera then captures the resultant surface point clouds of the materials as task goals (see Figure 4). We compare our skill optimisation method to direct trajectory optimisation (controlling 6D Cartesian displacement at every global step), Covariance Matrix Adaptation MAP-Annealing (CMA-MAE) [39, 40] and goal-conditioned SAC (see Appendix C for implementation details).
@@
CMA-MAE is a state-of-the-art evolutionary algorithm that does not use gradients to optimise a solution but optimises the problem by searching for a batch of high-quality and diverse solutions [39, 40]. For each task, it is run with two optimisation directions (i.e., two emitters) with a solution batch size of $10$ for $20$ iterations with a stepsize of $0.03$.
@@ VI RESULTS - SYSTEM IDENTIFICATION
This section presents the system identification task performance with differentiable physics and the ablation study that helps to determine the effectiveness of the EMD and HMD losses, the gradient operations (clipping, dynamic-scaling and normalisation), and the line search technique (five random seeds). To recall, we follow  [5] to conduct the DPSI task by optimising four physics parameters (Young’s modulus, Poisson’s ratio, material density, and sand friction angle) with gradients from either the EMD or HMD loss between the simulated and real manipulated material deformed by the same end-effector motion. In this work, we perform DPSI with soil and sand. As performing optimisation with a good initial solution accelerates the convergence of DPSI suggested by [5], we also present the performance of starting from initial parameter values based on [38] (denoted by “-ManInit”).
The results are presented in Figure 6. First of all, it can be seen that the optimisation diverges with both materials and loss functions (the red and pink lines) while the use of line search greatly stabilises gradient descent and converges to solutions with much reduced variances. This evidence demonstrates that our hypothesis of the optimisation being affected by a rugged loss landscape is accurate and the line search method is effective in addressing this problem. Secondly, with a good initial solution (dark and light green curves), DPSI can converge faster to better solutions with lower variances in both materials, aligned with [5]. Thirdly, the normalisation (dark and light blue curves) and dynamic-scaling (dark and light purple curves) operations perform worse than simply clipping the gradients (dark and light green curves). Lastly, the HMD loss performs slightly better than the EMD loss as all the light-coloured curves/scatter-points show slightly improved values than the dark-coloured ones (specifically, lower validation losses and variances).
@@ VIII-A Baselines
The benchmark result provides a few insights to guide research and engineering practice. First, optimising in the skill parameter space is significantly more efficient than in the trajectory space, even with demonstrations. Secondly, when an efficient and physically realistic simulator is available, learning a neural network policy with reinforcement learning methods was found to perform worse compared to gradient descent or MPC algorithms. This is somewhat unsurprising because a DRL method is essentially optimising a significantly larger parameter space, in contrast to the skill parameters of size five. It suggests that physically realistic simulators, preferably differentiable, are of higher potential and importance to efficient and high-quality robotic manipulation. Lastly, our gradient descent method and the model predictive control (CMA-MAE) method perform comparably well in terms of both solution quality and computational cost in our task setting. This is aligned with comparisons in other domains, such as image generation or classic RL tasks [41]. In practice, the hyperparameters of evolutionary algorithms, such as CMA-MAE, are generally more difficult to tune. The evolving direction of its stochastic search depends fully on the cost function, lacking consideration of the connection from the solution through the system dynamics to the post-manipulation states. This could lead to physically unreasonable solutions in complex tasks.
@@ IX-B Differentiability of the constitutive models
We employ the Saint Venant–Kirchhoff (SVK) model as our elastic energy density function (constitutive model) and the Drucker-Prager (DP) following [13]. In our work, we require the gradients of key physical properties. To verify the differentiability of this process, we manually calculate the partial derivatives of the input variables (i.e., the Lamé constants of the soil) with respect to the output of the elastoplasticity model (Eqs. 3 to 7). In the DP plasticity model, for cases without projection and projection onto the cone tip, we have $\frac{\partial S^{\prime}_{p}}{\partial\mu}=\frac{\partial S^{\prime}_{p}}{\partial\lambda}=O_{dim}$, where $O_{dim}$ denotes the zero matrix of dimension dimdim. In contrast, for projection onto the cone surface:
@@ IX-C Goal-conditioned soft actor-critic
where $\alpha_{sac}$ is a weight that controls the importance of maximising the policy entropy during policy optimisation, automatically updated during runtime following [43]. To reduce value overestimation, we maintain two q networks following [44] and take the minimal value in policy optimisation. The objective of both critic networks in this one-step POGCMDP is to minimise the following mean-square error:
@@
As indicated in [42], goal-relabelling helps to improve sample efficiency for goal-conditioned RL. Therefore, for every interaction experience ($o,\hat{g},\bar{a},r,o^{\prime}$) collected during training, we duplicate the experience, replace the desired goal with the achieved goal (i.e., the resultant observation $o^{\prime}\rightarrow\hat{g}$), recompute the goal-conditioned reward, and append the relabelled experience into the replay buffer.
Implementation details: To process point cloud data, we employ the PointNet network architecture for a shared encoder for all q and policy networks [45]. For the q networks, the encoded feature is passed to two MLP layers of size $1024$ and $512$ with batch normalisation and ReLU activation, with dropout ($0.4$) applied to the second layer, and then passed to the last MLP layer of size $256$, outputting the q value. For the policy network, the encoded feature is passed to three MLP layers of sizes $1024$, $256$ and $256$ with ReLU activation and then two separate MLP layers of size $256$ that output the mean (activated by the hyperbolic tangent function) and the logarithm of the standard deviation of a Gaussian policy distribution of the action’s dimension. The parameters of the PointNet encoder are only updated through the critic loss function, while the policy objective is used to update the MLP layers only.

### v4

@@ III-E Stopping gradients from exploding
A critical issue that destabilises or even disables gradient-based optimisation with differentiable physics is the numerical instability of gradient computation caused by long task horizons and the large particle number. As mentioned in existing works, gradients tend to explode or vanish as the length of the computation chain (number of actions/simulation steps) or the number of particles increases [undefg, undefa]. This effect can be predicted based on Eq. 35, where $\partial o_{t+1}/\partial a_{t}$ depends heavily on the number of simulation steps $T$, substeps $N$, and particles $|o_{t+1}|$. It increases drastically (explodes) when there are too many sub-gradients that are greater than 11 ($\partial f_{sub}/\partial\boldsymbol{x}_{p}>1$), and vice versa (vanishes). This problem is exacerbated when the physical model becomes more complex, preventing previous works from using efficient first-order optimisation methods on tasks with long horizons or large numbers of particles [undefg, undefh, undefa, undefb].
@@ IV-A Real-world platform
We set up a UR5e robotic arm to perform digging manipulations in the real world. As shown in Figure 3, the robot is equipped with a 3D-printed shovel as the end-effector. We place a $0.28\times 0.28$ m2 wooden box with $0.11$ metre depth to contain sand or soil for manipulation next to the robot. In the simulation, we use a box of the same size and the 3D model of the shovel as the simulated end-effector. We fill the container with the material, which could be sand or solid, to a height of $0.07$ metres. A Zivid Medium One+ high-resolution camera [undefaj] is used to collect real-world surface point clouds.
@@ IV-B Soil/sand system identification
Before running skill learning and parameter optimisation with the simulation dynamics, an important task is to ensure that the simulated soil/sand dynamics match the real ones as closely as possible. To achieve this, we follow the Differentiable Physics-based System Identification (DPSI) framework introduced in work [undefd], which can identify physics parameters with only one datapoint. In simple terms, the DPSI framework adjusts the key parameters of a physics model using gradients backpropagated from a loss function that compares the simulated and real point clouds of the same manipulation motion.
@@
The ranges of the physics parameters are defined according to an open-source structural engineering dataset [undefak]. They are summarised together with the gradient descent step sizes in Table II. The RMSprop algorithm with $\beta=0.9$ [undefn] is used for gradient-based optimisation in this work.
@@
TABLE II: Stepsizes and the ranges of values for the physics parameters: Young’s modulus $E$, Poisson’s ratio $\nu$, material density $\rho$, and sand friction angle $\phi_{f}$, selected based on an open-source structural engineering dataset [undefak].
The feasibility of DPSI for physics parameter identification has been demonstrated with playdough-like plastic objects, which are governed by the fixed-corotated elastic energy model and the von Mises plasticity model [undefd]. In this paper, we cope with granular materials that are modelled by the St. Venant-Kirchoff (SVK) elastic energy and the Drucker-Prager (DP) plastic yield criterion [undefl].
@@ IV-C Digging tasks
After the simulation is aligned with the real materials, we examine the feasibility of gradient-based skill optimisation. We manually use the shovel to create $3$ digging target shapes with both soil and sand. The optical camera then captures the resultant surface point clouds of the materials as task goals (see Figure 4). We compare our skill optimisation method to direct trajectory optimisation (controlling 6D Cartesian displacement at every global step), Covariance Matrix Adaptation MAP-Annealing (CMA-MAE) [undefal, undefam] and goal-conditioned SAC (see Appendix C for implementation details).
@@
CMA-MAE is a state-of-the-art evolutionary algorithm that does not use gradients to optimise a solution but optimises the problem by searching for a batch of high-quality and diverse solutions [undefal, undefam]. For each task, it is run with two optimisation directions (i.e., two emitters) with a solution batch size of $10$ for $20$ iterations with a stepsize of $0.03$.
@@ VI RESULTS - SYSTEM IDENTIFICATION
This section presents the system identification task performance with differentiable physics and the ablation study that helps to determine the effectiveness of the EMD and HMD losses, the gradient operations (clipping, dynamic-scaling and normalisation), and the line search technique (five random seeds). To recall, we follow  [undefd] to conduct the DPSI task by optimising four physics parameters (Young’s modulus, Poisson’s ratio, material density, and sand friction angle) with gradients from either the EMD or HMD loss between the simulated and real manipulated material deformed by the same end-effector motion. In this work, we perform DPSI with soil and sand. As performing optimisation with a good initial solution accelerates the convergence of DPSI suggested by [undefd], we also present the performance of starting from initial parameter values based on [undefak] (denoted by “-ManInit”).
The results are presented in Figure 6. First of all, it can be seen that the optimisation diverges with both materials and loss functions (the red and pink lines) while the use of line search greatly stabilises gradient descent and converges to solutions with much reduced variances. This evidence demonstrates that our hypothesis of the optimisation being affected by a rugged loss landscape is accurate and the line search method is effective in addressing this problem. Secondly, with a good initial solution (dark and light green curves), DPSI can converge faster to better solutions with lower variances in both materials, aligned with [undefd]. Thirdly, the normalisation (dark and light blue curves) and dynamic-scaling (dark and light purple curves) operations perform worse than simply clipping the gradients (dark and light green curves). Lastly, the HMD loss performs slightly better than the EMD loss as all the light-coloured curves/scatter-points show slightly improved values than the dark-coloured ones (specifically, lower validation losses and variances).
@@ VIII-A Baselines
The benchmark result provides a few insights to guide research and engineering practice. First, optimising in the skill parameter space is significantly more efficient than in the trajectory space, even with demonstrations. Secondly, when an efficient and physically realistic simulator is available, learning a neural network policy with reinforcement learning methods was found to perform worse compared to gradient descent or MPC algorithms. This is somewhat unsurprising because a DRL method is essentially optimising a significantly larger parameter space, in contrast to the skill parameters of size five. It suggests that physically realistic simulators, preferably differentiable, are of higher potential and importance to efficient and high-quality robotic manipulation. Lastly, our gradient descent method and the model predictive control (CMA-MAE) method perform comparably well in terms of both solution quality and computational cost in our task setting. This is aligned with comparisons in other domains, such as image generation or classic RL tasks [undefan]. In practice, the hyperparameters of evolutionary algorithms, such as CMA-MAE, are generally more difficult to tune. The evolving direction of its stochastic search depends fully on the cost function, lacking consideration of the connection from the solution through the system dynamics to the post-manipulation states. This could lead to physically unreasonable solutions in complex tasks.
@@ IX-B Differentiability of the constitutive models
We employ the Saint Venant–Kirchhoff (SVK) model as our elastic energy density function (constitutive model) and the Drucker-Prager (DP) following [undefl]. In our work, we require the gradients of key physical properties. To verify the differentiability of this process, we manually calculate the partial derivatives of the input variables (i.e., the Lamé constants of the soil) with respect to the output of the elastoplasticity model (Eqs. 3 to 7). In the DP plasticity model, for cases without projection and projection onto the cone tip, we have $\frac{\partial S^{\prime}_{p}}{\partial\mu}=\frac{\partial S^{\prime}_{p}}{\partial\lambda}=O_{dim}$, where $O_{dim}$ denotes the zero matrix of dimension dimdim. In contrast, for projection onto the cone surface:
@@ IX-C Goal-conditioned soft actor-critic
where $\alpha_{sac}$ is a weight that controls the importance of maximising the policy entropy during policy optimisation, automatically updated during runtime following [undefap]. To reduce value overestimation, we maintain two q networks following [undefaq] and take the minimal value in policy optimisation. The objective of both critic networks in this one-step POGCMDP is to minimise the following mean-square error:
@@
As indicated in [undefao], goal-relabelling helps to improve sample efficiency for goal-conditioned RL. Therefore, for every interaction experience ($o,\hat{g},\bar{a},r,o^{\prime}$) collected during training, we duplicate the experience, replace the desired goal with the achieved goal (i.e., the resultant observation $o^{\prime}\rightarrow\hat{g}$), recompute the goal-conditioned reward, and append the relabelled experience into the replay buffer.
Implementation details: To process point cloud data, we employ the PointNet network architecture for a shared encoder for all q and policy networks [undefar]. For the q networks, the encoded feature is passed to two MLP layers of size $1024$ and $512$ with batch normalisation and ReLU activation, with dropout ($0.4$) applied to the second layer, and then passed to the last MLP layer of size $256$, outputting the q value. For the policy network, the encoded feature is passed to three MLP layers of sizes $1024$, $256$ and $256$ with ReLU activation and then two separate MLP layers of size $256$ that output the mean (activated by the hyperbolic tangent function) and the logarithm of the standard deviation of a Gaussian policy distribution of the action’s dimension. The parameters of the PointNet encoder are only updated through the critic loss function, while the policy objective is used to update the MLP layers only.